# Context-Aware Pseudo-Label Scoring for Zero-Shot Video Summarization

### v1

@@ 2. Related Work
Video summarization has evolved from supervised sequence modeling to weaker-supervision and prompting paradigms with large language models (LLMs). Early studies relied on dense human annotations to learn frame-/shot-level importance; subsequent works introduced unsupervised and adversarial objectives to reduce labeling costs. Recently, prompting-based pipelines have enabled training-free (zero-shot) summarization and rubric-guided evaluation, but they remain sensitive to handcrafted templates and dataset-specific normalization. In parallel, pseudo labeling has re-emerged as a bridge between weak supervision and prompt-based reasoning, improving stability and semantic consistency. We review these lines with a focus on generic and query-focused settings.
Traditional video summarization. Methods are commonly grouped as supervised, weak/self-supervised, and unsupervised. Supervised models learn an importance function from dense annotations and cast summarization as sequence modeling to capture long-range dependencies, achieving strong results on SumMe and TVSum (Gygli et al., 2014a; Song et al., 2015a; Zhang et al., 2016a). Weak/self-supervised approaches reduce annotation demands by leveraging coarse cues (e.g., tags, user summaries, captions, or pseudo labels) and optimizing representativeness/diversity (Gygli et al., 2015). Unsupervised and adversarial/self-reconstruction methods (e.g., SUM-GAN) (Mahasseni et al., 2017) balance compactness and fidelity without labels, but often struggle to capture dataset-specific annotation styles and align with human semantics in a controllable, interpretable manner.
@@
Training-free prompting pipelines. Zero-shot systems describe scenes with vision–language models and employ LLMs to score importance via handcrafted templates, achieving competitive results on SumMe/TVSum (Barbara and Maalouf, 2025). However, they are highly sensitive to wording and template design and typically require dataset-specific score normalization (e.g., different transforms for SumMe vs. TVSum) (Barbara and Maalouf, 2025; Gygli et al., 2014a; Song et al., 2015a), revealing limited dataset adaptivity.
@@
Pseudo labeling for optimization and prompt stability. Pseudo labeling is a long-established self-training strategy that exploits unlabeled data via confidence-based supervision (Lee, 2013), and has been integrated with consistency regularization in Noisy Student (Xie et al., 2020) and FixMatch (Sohn et al., 2020) to improve efficiency and robustness. In LLMs, pseudo labeling extends to prompt/evaluation optimization: Self-Instruct (Wang et al., 2023) expands task templates with pseudo-instructions; RubricLLM (Hashemi and et al., 2024) and LLM-as-a-Judge (Zheng et al., 2023) use pseudo-derived rubrics to mitigate bias and enhance reliability; and pseudo labels assist prompt alignment and reward modeling for stable judgments (Zhou et al., 2024). In video summarization, pseudo labels promote semantic relevance and temporal coherence (e.g., AC-SUM-GAN (Apostolidis et al., 2020), Cycle-SUM/RS-SUM (Yuan et al., 2019)).
Query-focused video summarization (QFVS). Query-focused summarization explicitly conditions the summary on a user query. SH-DPP introduced a sequential–hierarchical DPP to jointly model query relevance and contextual importance, establishing the query-focused setting (Sharghi et al., 2016a). The QFVS benchmark further provided a dataset, evaluation protocol, and a memory-network parameterized sequential DPP to align textual queries with temporal shots (Sharghi et al., 2017). More recently, CLIP-It unified generic and query-focused summarization under a language-guided multimodal transformer, delivering strong supervised performance and becoming a common reference baseline (Narasimhan et al., 2021). While these methods advance the modeling of query–video alignment and the relevance–diversity trade-off, template sensitivity and limited dataset adaptivity remain open issues—motivating rubric-guided, pseudo-labeled prompting with contextualized scoring as pursued in this work.
Figure 2. Overall architecture of the proposed framework.
@@ 3.1. Scenario Division
To divide the video into consecutive segments, we use two stages. In the first stage, we used a preliminary vision-based segmentation, and in the second stage, we used a semantic-based boundary enhancement strategy. Partition the input video into structurally coherent segments (shots) so that each segment corresponds to a relatively homogeneous visual context. This segmentation provides the basic units for subsequent summarization, ranking, and selection.
Step 1. Frame preprocessing. Given a video frame $I_{t}$, it is first converted to grayscale and resized to a fixed resolution $N\times N$:
@@
Step 2. Frequency transform (DCT). A two-dimensional discrete cosine transform (DCT) is applied to $X_{t}$ to obtain frequency coefficients:
(2) $C_{t}=\mathrm{DCT2}(X_{t}).$
The low-frequency components preserve the global structural information, while high-frequency coefficients mainly capture texture and noise.
Step 3. Low-frequency block and mean. The top-left $k\times k$ block of $C_{t}$ is extracted as the low-frequency region:
(3) $A_{t}=C_{t}[0:k,\,0:k].$
Excluding the DC term $A_{t}(0,0)$, the mean of the remaining coefficients is computed as:
(4) $\mu_{t}=\frac{1}{k^{2}-1}\sum_{\begin{subarray}{c}0\leq i,j<k\\ (i,j)\neq(0,0)\end{subarray}}A_{t}(i,j).$
Step 4. Hash generation. Each coefficient in $A_{t}$ is compared with the mean value $\mu_{t}$, yielding a binary hash:
(5) $h_{t}^{(i,j)}=\begin{cases}1,&A_{t}(i,j)>\mu_{t},\\ 0,&A_{t}(i,j)\leq\mu_{t},\end{cases}\qquad(i,j)\neq(0,0).$
Flattening these values produces a binary vector $h_{t}\in\{0,1\}^{M}$ with $M=k^{2}-1$.
@@
Step 5. Frame-wise difference (normalized Hamming distance). The structural difference between two consecutive frames is defined as:
(6) $D(t,t+1)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{1}\!\left[h_{t}^{(m)}\neq h_{t+1}^{(m)}\right],$
where $\mathbb{1}[\cdot]$ is the indicator function.
@@
Step 6. Scene boundary decision. A scene boundary is declared when the normalized Hamming distance exceeds a threshold $\tau$:
(7) $\text{Boundary}(t)=\begin{cases}1,&D(t,t+1)\geq\tau,\\ 0,&\text{otherwise}.\end{cases}$
Step 7. Adaptive Threshold Selection. Instead of relying on a fixed threshold $\tau$ to detect scene boundaries, we employ a dynamic strategy to automatically determine the optimal threshold $\tau^{\star}$. A set of candidate values is defined as
$\{\tau_{\min},\tau_{\min}+\Delta\tau,\ldots,\tau_{\max}\},$
and for each candidate $\tau_{i}$, we compute the number of detected scenes $N(\tau_{i})$. The rate of change in scene count with respect to the threshold is then approximated by a finite difference:
(8) $\Delta N(\tau_{i})=\frac{N(\tau_{i}+\Delta\tau)-N(\tau_{i})}{\Delta\tau}.$
The optimal threshold is selected as the value that maximizes the magnitude of the negative slope:
(9) $\tau^{\star}=\arg\max_{\tau_{i}}\;-\,\Delta N(\tau_{i}).$
In practice, the number of detected scenes $N(\tau)$ typically increases gradually at low threshold values, reaches a peak, and then drops sharply as $\tau$ continues to increase. The chosen threshold $\tau^{\star}$ corresponds to the point of maximum negative slope after the peak, which captures the most significant reduction in scene counts and thus provides a stable operating point for segmentation.
@@
Step 8. Scene Boundary Refinement. To enhance temporal coherence and semantic consistency, we refine the initial scene segmentation by eliminating overly short segments (Diao et al., 2025; Yao and Chuah, 2025). Such short scenes often arise from noise or subtle visual fluctuations and typically lack sufficient content, leading to Scene Boundary Refinement. To enhance temporal coherence and semantic consistency, we refine the initial scene segmentation by eliminating overly short segments. Such short scenes often arise from noise or subtle visual fluctuations and typically lack sufficient content, leading to unnecessary fragmentation (Wu et al., 2025a, b). We regard a scene as short if its length is less than $M=150$ frames(approximately 5 seconds). In practice, we observe that setting a threshold of 5 seconds helps prevent over-segmentation without removing meaningful events. The initial segmentation stage identifies boundaries primarily through sharp changes in visual features. To further refine these boundaries, we merge back any short segment whose visual representation is highly similar to one of its neighboring segments, ensuring smoother and more consistent temporal units. For a given short scene $S_{i}$, we first compute its mean embedding:
@@ 3.4. Scene Scoring Prompt Design
We guide the LLM to assign scene-level importance scores through carefully engineered prompts grounded in data rather than ad-hoc templates. First, to construct reliable pseudo labels for prompt design, we perform a reason-mining step: for each video we select the top-3 highest-scoring and top-3 lowest-scoring segments (obtained from the GT-derived scene scores), feed their captions into GPT-4o with a strict-JSON prompt, and elicit concise explanations of why the high-score scenes matter, why the low-score scenes do not, and what discriminates the two. We aggregate these rationales into a structured, dataset-adaptive rubric that specifies evaluation dimensions and penalties, then instantiate two rubric-guided scoring prompts tailored to TVSum and SumMe.
@@
Once representative high- and low-score segments are identified, we employ the LLM to perform reason mining. To ensure balanced and representative supervision, we select the top-$3$ highest-scoring and top-$3$ lowest-scoring segments from each video as reference samples, based on the ground-truth frame-level scores. The textual descriptions of these selected segments—derived from the scene division stage—are fed into GPT-4o with a structured prompt that enforces strict JSON-format outputs. The model is required to summarize the distinguishing characteristics between high- and low-importance scenes and articulate the underlying reasons that contribute to their scores. These extracted rationales are then distilled into dataset-specific rubric blocks, which serve as the foundation for constructing our final scoring prompts (see Appendix A), while the core part is shown below. The JSON object contains three fields:
@@
After collecting reasons for high- and low-score segments, we move beyond manual inspection by introducing a second-stage abstraction prompt that compels the LLM to consolidate these cues into a hierarchical, dataset-aware rationale schema. Concretely, the LLM is instructed to: (i) cluster recurring positive/negative/difference cues across videos; (ii) elevate them into evaluation dimensions with explicit constraints and penalties; (iii) add category-specific checklists capturing dataset idiosyncrasies; and (iv) formalize a calibration ladder and an exact $[0,100]$ output rule. The result is a multi-level, interpretable rationale chain—evidence $\rightarrow$ cue $\rightarrow$ dimension $\rightarrow$ scoring rule—that preserves dataset characteristics while remaining generalizable. This structured rationale is then compiled into rubric-guided scoring prompts used by GPT-5 at inference, enabling consistent, bias-resistant scene scoring and reducing sensitivity to prompt wording. TVSum rubric prompt. The complete prompt is provided in Appendix D. Below shows the main scoring rubric for TVSum:
@@
SumMe rubric prompt. The complete prompt is provided in Appendix C. Below summarizes the major scoring rules for SumMe:
1) Title/Thematic Relevance (0–35)
- Directly reflects the video title’s subject (e.g., aircraft motion, bus in tunnel, chef cooking).
- Low if off-focus (e.g., generic close-ups, irrelevant context).
2) Action/Interaction & Skill (0–20)
- Visible motion, operation, or behavior (landing, jumping, teamwork).
- Static or idle scenes are low unless critical setup.
3) Detail & Visibility (0–15)
- Informative clarity: tools, textures, instruments, or gestures.
- Penalize blur, darkness, occlusion.
4) Informational Uniqueness (0–15)
- Adds new step or viewpoint; penalize near-duplicates.
5) Narrative Progression (0–15)
- Shows setup→action→result or transformation moment.
Penalties: off-topic (-10), static (-8), low visibility (-6), redundancy (-6).
Category checklists: apply domain-specific cues (aviation, cooking, sports, etc.).
Calibration ladder: ensures score spread and penalizes inflation.
Final score = round(0.35R + 0.20A + 0.15D + 0.15U + 0.15N - penalties), clamp [0,100].
Output exactly one integer in [0,100].
QFVS rubric prompt. The complete prompt is provided in Appendix E. Below summarizes the major scoring rules for QFVS:
1) Destination / Task Anchoring (0--30)
- Strong: clear venue or task anchors (e.g., restaurant interior, supermarket aisle, classroom, workstation).
- Weak: generic street or decor shots lacking contextual link.
2) Action $\rightarrow$ Visible Outcome (0--20)
- Strong: concrete, goal-oriented interactions with visible results (e.g., ordering, paying, teaching, operating).
- Weak: idle presence or repetitive motion without new effect.
3) State Change & Narrative Progress (0--20)
[…]
- Weak: no progression or unrelated cutaways.
4) Evidence & Identifiers (0--15)
- Strong: readable cues verifying context (menus, receipts, books, labels, screens).
- Weak: partial or illegible details, unverified hints.
5) Visual Clarity & Focus (0--15)
- Strong: key actors and objects unobstructed, well-lit, and interpretable.
- Penalty for dark, blurred, or occluded frames.
Penalties: off-topic cutaway (-10), static ambience (-8), redundancy (-6), low visibility or indistinct text (-6).
[…]
Calibration ladder: 90--100 (essential action), 75--89 (strong task), 60--74 (supportive context), 40--59 (weak linkage), 0--39 (off-topic).
Query modifier: $\pm$(0--10) on Dimension 1 for strong/poor alignment with query theme.
Final score = round(0.30A + 0.20C + 0.20S + 0.15E + 0.15V + modifiers), clamp [0,100].
Output one integer in [0,100] per query.
@@ 3.4.3. Score generation with contextualized inputs
We use two prompts; the full templates are in Appendix D. Below are the operational instructions actually fed to the LLM.
@@ 4.4. Ablation Study
Refer to caption Refer to caption Refer to caption
(a) F1 (b) Precision (c) Recall
Figure 3. Bar charts on TVSum and SumMe. For each dataset, the three bars correspond respectively to Baseline (yellow), +Pseudo Label (green), and +Pseudo +Context (blue). Each metric (F1, Precision, Recall) shows a consistent improvement as pseudo-label reconstruction and contextual prompting are applied.
@@
This study investigates two components of our zero-shot framework: (i) pseudo-label reconstruction prompts and (ii) contextualized scoring, where the reconstructed prompts are augmented with local scene context (summaries of the preceding and following segments). All variants use identical data splits and evaluation protocols so that performance differences arise solely from prompt design. The quantitative results on SumMe/TVSum (Table 4) together with the bar-plot visualizations (Figure 3) and the QFVS ablations (Table 5, cf. Table 3) jointly substantiate the effectiveness of the proposed framework.
Effects on SumMe/TVSum (Table 4; Figure 3). Relative to the plain zero-shot baseline, introducing pseudo-label reconstruction increases F1 from 56.73 to 57.24 on SumMe (+0.51) and from 62.21 to 62.83 on TVSum (+0.62), accompanied by consistent gains in precision and recall. Adding contextualized scoring on top of pseudo labels further raises F1 to 57.58 on SumMe (+0.34 over pseudo labels; +0.85 over baseline) and to 63.05 on TVSum (+0.22 over pseudo labels; +0.84 over baseline). These results indicate that short summaries of adjacent scenes enhance temporal coherence and reduce redundancy, yielding more consistent frame-importance distributions.
@@ Acknowledgment
We acknowledge the use of GPT-4o (OpenAI, 2024) for grammar correction, typo fixing, and partial sentence rephrasing to improve the clarity and readability of this paper.
Appendix A Reason Mining Prompt
Appendix B TVSum Rubric Prompt
Appendix C SumMe Rubric Prompt
The detailed SumMe Rubric Prompt, including dataset-specific category checklists, can be found at: https://github.com/wuyuanli60-svg/Context-Aware-Pseudo-Label-Scoring-for-Zero-Shot-Video-Summarization/blob/main/Appendix_C_SumMe_Rubric_Prompt.txt
Appendix D Context-Aware Scoring Prompts
Appendix E Context-Aware Scoring Prompts

### v3

@@ 2. Related Work
Video summarization has progressed from supervised sequence models to weaker supervision and prompt-based pipelines with large language models (LLMs). Early work learned frame/shot importance from dense labels and long-range dependencies, showing strong results on SumMe/TVSum (Gygli et al., 2014a; Song et al., 2015a; Zhang et al., 2016a); later methods reduced annotation with weak/self-supervision and submodular/unsupervised objectives (Gygli et al., 2015; Mahasseni et al., 2017).
@@
Training-free prompting. Recent zero-shot systems describe scenes with vision–language models and rely on LLM prompts to score importance, reaching competitive performance but remaining sensitive to template wording and dataset-specific normalization (Barbara and Maalouf, 2025; Gygli et al., 2014a; Song et al., 2015a).
@@
Pseudo labeling. Pseudo labels (self-training) improve stability and semantic consistency, historically via confidence- or consistency-based supervision (Lee, 2013; Xie et al., 2020; Sohn et al., 2020), and more recently for prompt/evaluation optimization—e.g., expanding task instructions (Wang et al., 2023), rubric-guided judging (Zheng et al., 2023; Hashemi and et al., 2024), and alignment for reliable scoring (Zhou et al., 2024). In video summarization, pseudo labels encourage relevance and temporal coherence (e.g., AC-SUM-GAN, Cycle-SUM/RS-SUM) (Apostolidis et al., 2020; Yuan et al., 2019).
Query-focused video summarization (QFVS). QFVS conditions summaries on a user query; SH-DPP and the QFVS benchmark established the setting and protocols (Sharghi et al., 2016a, 2017), while CLIP-It unified generic and query-focused summarization under language guidance (Narasimhan et al., 2021). Remaining challenges include template sensitivity and dataset adaptivity, motivating our rubric-guided, pseudo-labeled prompting with contextualized scoring.
Chain-of-thought (CoT) optimization. CoT prompting and its variants provide a practical way to improve LLM reasoning reliability—few/zero-shot CoT (Wei et al., 2022; Kojima et al., 2022) and self-consistency decoding (Wang et al., 2022) reduce prompt brittleness; for evaluation, LLM-as-a-Judge and rubric-based judging improve consistency and calibration (Zheng et al., 2023; Hashemi and et al., 2024). Our design leverages these insights to stabilize rubric-guided scoring while retaining zero-shot generality.
@@ 3.1. Scenario Division
We split each video into consecutive, semantically coherent scenes with a lightweight vision-based detector and an adaptive threshold. Perceptual hashes are computed using the Python ImageHash library (imagehash.phash, PIL backend).
Step 1. Frame preprocessing. Given a frame $I_{t}$, convert to grayscale and resize to a fixed resolution $N\times N$:
@@
Step 2. Perceptual hash (pHash). Apply imagehash.phash to $X_{t}$ to obtain a binary hash vector $h_{t}\in\{0,1\}^{M_{h}}$.
@@
Step 3. Frame-wise difference (normalized Hamming). The structural change between consecutive frames is
(2) $D(t,t{+}1)\;=\;\frac{1}{M_{h}}\sum_{m=1}^{M_{h}}\mathbb{1}\!\left[h_{t}^{(m)}\neq h_{t+1}^{(m)}\right].$
Symbols: $M_{h}$ is hash length (bits); $h_{t}^{(m)}$ is the $m$-th bit of $h_{t}$; $\mathbb{1}[\cdot]$ is the indicator.
@@
Step 4. Scene boundary decision. Declare a boundary when the change exceeds a threshold:
(3) $\mathrm{Boundary}(t)=\mathbb{1}\!\left[D(t,t{+}1)\geq\tau\right].$
Step 5. Adaptive threshold selection. Instead of a fixed $\tau$, select $\tau^{\star}$ on a grid $\{\tau_{\min},\tau_{\min}{+}\Delta\tau,\ldots,\tau_{\max}\}$ using the steepest post-peak drop of the scene-count curve:
(4) $\Delta N(\tau_{i})\;=\;\frac{N(\tau_{i}{+}\Delta\tau)-N(\tau_{i})}{\Delta\tau},$
(5) $\tau^{\star}\;=\;\arg\max_{\tau_{i}}\big(-\,\Delta N(\tau_{i})\big).$
Symbols: $N(\tau)$—number of detected scenes at threshold $\tau$; $\Delta\tau$—grid step; $\tau^{\star}$—selected operating threshold.
@@
Step 6. Scene Boundary Refinement. To enhance temporal coherence and semantic consistency, we refine the initial scene segmentation by eliminating overly short segments (Diao et al., 2025; Yao and Chuah, 2025). Such short scenes often arise from noise or subtle visual fluctuations and typically lack sufficient content, leading to Scene Boundary Refinement. To enhance temporal coherence and semantic consistency, we refine the initial scene segmentation by eliminating overly short segments. Such short scenes often arise from noise or subtle visual fluctuations and typically lack sufficient content, leading to unnecessary fragmentation (Wu et al., 2025a, b). We regard a scene as short if its length is less than $M=150$ frames(approximately 5 seconds). In practice, we observe that setting a threshold of 5 seconds helps prevent over-segmentation without removing meaningful events. The initial segmentation stage identifies boundaries primarily through sharp changes in visual features. To further refine these boundaries, we merge back any short segment whose visual representation is highly similar to one of its neighboring segments, ensuring smoother and more consistent temporal units. For a given short scene $S_{i}$, we first compute its mean embedding:
@@ 3.4. Scene Scoring Prompt Design
We guide the LLM to assign scene-level importance scores through carefully engineered prompts grounded in data rather than ad-hoc templates. First, to construct reliable pseudo labels for prompt design, we perform a reason-mining step: for each video we select the top-3 highest-scoring and top-3 lowest-scoring segments (obtained from the GT-derived scene scores), feed their captions into GPT-4o with a strict-JSON prompt, and elicit concise explanations of why the high-score scenes matter, why the low-score scenes do not, and what discriminates the two. We aggregate these rationales into a structured, dataset-adaptive rubric that specifies evaluation dimensions and penalties, then instantiate rubric-guided scoring prompts tailored to TVSum (main text) and to SumMe/QFVS. All complete prompt designs (reason-mining prompt, dataset-specific rubric prompts, and contextual scoring templates) are provided in our repository: https://github.com/wuyuanli60-svg/Context-Aware-Pseudo-Label-Scoring-for-Zero-Shot-Video-Summarization.
@@
Once representative high- and low-score segments are identified, we employ the LLM to perform reason mining. To ensure balanced and representative supervision, we select the top-$3$ highest-scoring and top-$3$ lowest-scoring segments from each video as reference samples, based on the ground-truth frame-level scores. The textual descriptions of these selected segments—derived from the scene division stage—are fed into GPT-4o with a structured prompt that enforces strict JSON-format outputs. The model is required to summarize the distinguishing characteristics between high- and low-importance scenes and articulate the underlying reasons that contribute to their scores. These extracted rationales are then distilled into dataset-specific rubric blocks, which serve as the foundation for constructing our final scoring prompts, while the core part is shown below:
@@
After collecting reasons for high- and low-score segments, we move beyond manual inspection by introducing a second-stage abstraction prompt that compels the LLM to consolidate these cues into a hierarchical, dataset-aware rationale schema. Concretely, the LLM is instructed to: (i) cluster recurring positive/negative/difference cues across videos; (ii) elevate them into evaluation dimensions with explicit constraints and penalties; (iii) add category-specific checklists capturing dataset idiosyncrasies; and (iv) formalize a calibration ladder and an exact $[0,100]$ output rule. The result is a multi-level, interpretable rationale chain—evidence $\rightarrow$ cue $\rightarrow$ dimension $\rightarrow$ scoring rule—that preserves dataset characteristics while remaining generalizable. This structured rationale is then compiled into rubric-guided scoring prompts used by GPT-5 at inference, enabling consistent, bias-resistant scene scoring and reducing sensitivity to prompt wording.
TVSum rubric prompt (main text). Below shows the main scoring rubric for TVSum:
@@ 3.4.3. Score generation with contextualized inputs
We use two prompts. Below are the operational instructions actually fed to the LLM.
@@ 4.3. Query-Focused Video Summarization
4.4. Ablation Study
This study investigates two components of our zero-shot framework: (i) pseudo-label reconstruction prompts and (ii) contextualized scoring, where the reconstructed prompts are augmented with local scene context (summaries of the preceding and following segments). All variants use identical data splits and evaluation protocols so that performance differences arise solely from prompt design. The quantitative results on SumMe/TVSum (Table 4) and the QFVS ablations (Table 5) jointly substantiate the effectiveness of the proposed framework.
Effects on SumMe/TVSum (Table 4). Relative to the plain zero-shot baseline, introducing pseudo-label reconstruction increases F1 from 56.73 to 57.24 on SumMe (+0.51) and from 62.21 to 62.83 on TVSum (+0.62), accompanied by consistent gains in precision and recall. Adding contextualized scoring on top of pseudo labels further raises F1 to 57.58 on SumMe (+0.34 over pseudo labels; +0.85 over baseline) and to 63.05 on TVSum (+0.22 over pseudo labels; +0.84 over baseline). These results indicate that short summaries of adjacent scenes enhance temporal coherence and reduce redundancy, yielding more consistent frame-importance distributions.
@@ Acknowledgment
We acknowledge the use of GPT-5 (OpenAI, 2025) for grammar correction, typo fixing, and partial sentence rephrasing to improve the clarity and readability of this paper.